Complete { (2, 0, 3)} to form an orthogonal 

basis of R³


1
Expert's answer
2021-01-11T10:05:16-0500

Denote "u=(2,0,3)". A vector "v=(v_1,v_2,v_3)\perp u" if

"u.v=0\implies2v_1+3v_3=0\implies v_3=-\tfrac{2}{3}v_1\;(\text{and } v_2 \text{ is arbitrary})."

Take, for example "v_1=3,\;v_3=-2\text{ and }v_2=0". Then "v=(3,0,-2)" is orthogonal to "u".


To construct a third vector "w=(w_1,w_2,w_3)" orthogonal to "u\text{ and }v" we follow the same technique: "u.w\text{ and }v.w=0" give

"\begin{cases}2w_1+3w_3=0\\3w_1-2w_3=0\end{cases}\implies w_1=w_3=0\text{ and }w_2\text{ is arbitrary}"

Take, for example, "w=(0,1,0)".


Then "\{(2,0,3),(-3,0,2),(0,1,0)\}" is an orthogonal basis of "\mathbb{R}^3".